# Adaptive Gradient Normalization and Independent Sampling for (Stochastic) Generalized-Smooth Optimization

### v3

@@ Appendix E Proof of Theorem 1
where $(i)$ is using mean value theorem such that $\Phi(\Delta_{t+1})-\Phi(\Delta_{t})=|\Delta_{t+1}-\Delta_{t}\|\Phi^{\prime}(\xi)|$ and $\Delta_{t+1}^{-\theta}\geq|\Phi^{\prime}(\xi)|\geq\Delta^{-\theta}_{t}$ holds for any $\xi\in[\Delta_{t+1},\Delta_{t}]$; (ii) uses the fact $\Delta_{t+1}^{-\theta}\leq 2\Delta_{t}^{-\theta}$; (iii) utilizes recursion $\Delta_{t}-\Delta_{t+1}\geq\frac{\gamma(2\mu)^{\theta}}{4}\Delta_{t}^{\theta}$; (iv) uses the fact that $\Delta_{t}^{\theta}>\Delta_{t+1}^{\theta}$ for all $\theta>0$.
@@
When $\Delta_{t}$ is small enough, $\Delta_{t+1}^{{1}/{\omega}}$ will dominate $\Delta_{t+1}$ order-wisely since $11/\omega<$ 1. This fact leads to
@@ Appendix F Proof of Proposition 1
where (i) is due to descent lemma equation 1; (ii) and (iii) are due to the learning rate design $\gamma\leq\min\{\frac{1}{L_{0}},\frac{1}{2L_{1}G}\}$, which ensures $-(\gamma-\frac{L_{0}\gamma^{2}}{2})\leq-\frac{\gamma}{2}$ and $\frac{L_{1}\gamma^{2}}{2}\|\nabla f(w_{t})\|^{3}\leq\frac{\gamma}{4}\|\nabla f(w_{t})\|^{2}$. When applying Assumption 2 with $\rho=1$ and denote $f(w_{t})-f^{*}$ as $\Delta_{t}$, we have
@@ Appendix H Lemma 5 and Proof
<$\overset{(ii)}{<}$ $\frac{\gamma^{2}}{2h_{t}^{\beta}}\frac{L_{1}\big{\|}\nabla F(w_{t})\big{\|}^{\alpha}}{(4L_{1}\gamma)(\frac{1}{1-\tau_{1}}\big{\|}\nabla f_{\xi^{\prime}}(w_{t})\big{\|}+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\cdot\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
<$\overset{(iv)}{<}$ $\frac{\gamma^{2}}{2h_{t}^{\beta}}\frac{L_{1}}{4L_{1}\gamma}\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
where the first inequality uses the fact that $\frac{1}{h_{t}^{\beta}}\leq 1$; the second inequality uses the fact $\gamma\leq\frac{1}{4L_{0}(2\tau_{1}^{2}+1)}$; the third inequality uses the fact $\frac{4\tau_{2}^{2}}{2\tau_{1}^{2}+1}\leq 1+\frac{4\tau_{2}^{2}}{2\tau_{1}^{2}+1}$; the forth inequality uses the fact that $\big{\|}\nabla F(w_{t})\big{\|}^{2}>1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}}+1)$. Combining equation 49, equation 50, equation 51 gives us the desired result.
@@ Appendix I Proof of Theorem 2
Taking expectation over $\xi$,$\xi^{\prime}$ and $w_{t}$, and by the independence between $\xi$ and $\xi^{\prime}$, we have
@@
=$\overset{(i)}{=}$ $\gamma\min\Big{\{}1,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\gamma(\frac{1}{1-\tau_{1}}\big{\|}\nabla f_{\xi^{\prime}}(w_{t})\big{\|}+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\Big{\}}\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
=$\overset{(iv)}{=}$ $\min\Big{\{}\gamma,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\big{(}2C_{1}\big{\|}\nabla F(w_{t})\big{\|}\big{)}^{\beta}},\frac{1}{4L_{1}(2\tau_{1}^{2}+1)(2C_{2})^{\beta}}\Big{\}}\big{\|}\nabla F(w_{t})\big{\|}^{2}$
=$\overset{(v)}{=}$ $\min\Big{\{}\gamma,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\big{(}2C_{1}\big{\|}\nabla F(w_{t})\big{\|}\big{)}^{\beta}}\Big{\}}\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
where (i) expands the expression of $\frac{1}{h_{t}^{\beta}}$; (ii) utilizes the second inequality of Lemma 2 to upper bound $\|\nabla f_{\xi^{\prime}}(w_{t})\|$ by $(\tau_{1}+1)\|\nabla F(w_{t})\|+\tau_{2}$; (iii) utilizes the fact $\frac{1}{(a+b)^{\beta}}\geq\min\{\frac{1}{(2a)^{\beta}},\frac{1}{(2b)^{\beta}}\}$ by setting $a=\frac{\tau_{1}+1}{1-\tau_{1}}\big{\|}\nabla F(w_{t})\big{\|}$, $b=\frac{2\tau_{2}}{1-\tau_{1}}$, $\beta\geq 0$; (iv) substitute $\gamma$ inside the minimum operator. Moreover, (v) follows from the step size rule $\gamma=\frac{1}{8L_{1}(2\tau_{1}^{2}+1)(2\tau_{2}/(1-\tau_{1}))^{\beta}}\leq\frac{1}{4L_{1}(2\tau_{1}^{2}+1)(4\tau_{2}/(1-\tau_{1}))^{\beta}}$. Substituting equation 59 into equation 58 and summing over $t$ from 00 to $T-1$, we obtain that
@@
where (i) uses the concavity $\psi(\cdot)=y^{\frac{1}{2}}$ and $\psi(\cdot)=y^{\frac{1}{2-\beta}}$ and Jensen’s inequality $\psi(\frac{\sum_{i}x_{i}}{n})\geq\frac{\sum_{i}\psi(x_{i})}{n}$ applied on $\|\nabla F(w_{t})\|$; (ii) uses the fact $\mathbb{E}[\min\{A,B\}]\leq\mathbb{E}[A]\leq\max\{\mathbb{E}[A],\mathbb{E}[B]\}$; (iii) applies Jensen’s inequality over concave function $\Psi(\cdot)=y^{\frac{1}{2}}$ and $\Psi(\cdot)=y^{\frac{1}{2-\beta}}$; (iv) is due to re-organize equation 60, equation 61 and set $K={T}$, $K^{c}=\frac{T}{2}$ (Since we require either $K$ or $K^{c}$ should be larger than $\frac{T}{2}$ to guarantee that $K+K_{c}=T$).
@@
=$\overset{(i)}{=}$ $\gamma\min\Big{\{}1,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\gamma(\frac{1}{1-\tau_{1}}\big{\|}\nabla f_{\xi^{\prime}}(w_{t})\big{\|}+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\Big{\}}\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
=$\overset{(iv)}{=}$ $\min\Big{\{}\gamma,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\big{(}2C_{1}\big{\|}\nabla F(w_{t})\big{\|}\big{)}^{\beta}},\frac{1}{4L_{1}(2\tau_{1}^{2}+1)(2C_{2})^{\beta}}\Big{\}}\big{\|}\nabla F(w_{t})\big{\|}^{2}$
=$\overset{(v)}{=}$ $\min\Big{\{}\gamma,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\big{(}2C_{1}\big{\|}\nabla F(w_{t})\big{\|}\big{)}^{\beta}}\Big{\}}\big{\|}\nabla F(w_{t})\big{\|}^{2}$

### v4

@@ Appendix E Proof of Theorem 1
where $(i)$ applies mean value theorem such that $\Phi(\Delta_{t+1})-\Phi(\Delta_{t})=|\Delta_{t+1}-\Delta_{t}\|\Phi^{\prime}(\xi)|$ and $\Delta_{t+1}^{-\theta}\geq|\Phi^{\prime}(\xi)|\geq\Delta^{-\theta}_{t}$ holds for any $\xi\in[\Delta_{t+1},\Delta_{t}]$; (ii) uses the fact $\Delta_{t+1}^{-\theta}\leq 2\Delta_{t}^{-\theta}$; (iii) utilizes recursion $\Delta_{t}-\Delta_{t+1}\geq\frac{\gamma(2\mu)^{\theta}}{4}\Delta_{t}^{\theta}$; (iv) uses the fact that $\Delta_{t}^{\theta}>\Delta_{t+1}^{\theta}$ for all $\theta>0$.
@@
When $\Delta_{t}$ is small enough, $\Delta_{t+1}^{{1}/{\omega}}$ will dominate $\Delta_{t+1}$ order-wisely since $1/\omega<1$. This fact leads to
@@ Appendix F Proof of Proposition 1
where (i) is due to descent lemma equation 1; (ii) and (iii) are due to the learning rate design $\gamma\leq\min\{\frac{1}{L_{0}},\frac{1}{2L_{1}G}\}$, which ensures $(\gamma-\frac{L_{0}\gamma^{2}}{2})\leq-\frac{\gamma}{2}$ and $\frac{L_{1}\gamma^{2}}{2}\|\nabla f(w_{t})\|^{3}\leq\frac{\gamma}{4}\|\nabla f(w_{t})\|^{2}$.
When applying Assumption 2 with $\rho=1$ and denote $f(w_{t})-f^{*}$ as $\Delta_{t}$, we have
@@ Appendix H Lemma 5 and Proof
$\overset{(ii)}{<}$ $\frac{\gamma^{2}}{2h_{t}^{\beta}}\frac{L_{1}\big\|\nabla F(w_{t})\big\|^{\alpha}}{(4L_{1}\gamma)(\frac{1}{1-\tau_{1}}\big\|\nabla f_{\xi^{\prime}}(w_{t})\big\|+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\cdot\big\|\nabla F(w_{t})\big\|^{2}$
@@
$\overset{(iv)}{<}$ $\frac{\gamma^{2}}{2h_{t}^{\beta}}\frac{L_{1}}{4L_{1}\gamma}\big\|\nabla F(w_{t})\big\|^{2}$
@@
where the first inequality uses the fact that $\frac{1}{h_{t}^{\beta}}\leq 1$; the second inequality uses the fact $\gamma\leq\frac{1}{4L_{0}(2\tau_{1}^{2}+1)}$; the third inequality uses the fact $\frac{4\tau_{2}^{2}}{2\tau_{1}^{2}+1}\leq 1+\frac{4\tau_{2}^{2}}{2\tau_{1}^{2}+1}$; the fourth inequality uses the fact that $\big\|\nabla F(w_{t})\big\|^{2}>1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}}+1)$. Combining equation 49, equation 50, equation 51 gives us the desired result.
@@ Appendix I Proof of Theorem 2
Taking expectation over $\xi$$\xi^{\prime}$ and $w_{t}$, and by the independence between $\xi$ and $\xi^{\prime}$, we have
@@
$\overset{(i)}{=}$ $\gamma\min\Big\{1,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\gamma(\frac{1}{1-\tau_{1}}\big\|\nabla f_{\xi^{\prime}}(w_{t})\big\|+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\Big\}\big\|\nabla F(w_{t})\big\|^{2}$
@@
$\overset{(iv)}{=}$ $\min\Big\{\gamma,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\big(2C_{1}\big\|\nabla F(w_{t})\big\|\big)^{\beta}},\frac{1}{4L_{1}(2\tau_{1}^{2}+1)(2C_{2})^{\beta}}\Big\}\big\|\nabla F(w_{t})\big\|^{2}$
$\overset{(v)}{=}$ $\min\Big\{\gamma,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\big(2C_{1}\big\|\nabla F(w_{t})\big\|\big)^{\beta}}\Big\}\big\|\nabla F(w_{t})\big\|^{2}$
@@
where (i) expands the expression of $\frac{1}{h_{t}^{\beta}}$; (ii) utilizes the second inequality of Lemma 2 to upper bound $\|\nabla f_{\xi^{\prime}}(w_{t})\|$ by $(\tau_{1}+1)\|\nabla F(w_{t})\|+\tau_{2}$; (iii) utilizes the fact $\frac{1}{(a+b)^{\beta}}\geq\min\{\frac{1}{(2a)^{\beta}},\frac{1}{(2b)^{\beta}}\}$ by setting $a=\frac{\tau_{1}+1}{1-\tau_{1}}\big\|\nabla F(w_{t})\big\|$, $b=\frac{2\tau_{2}}{1-\tau_{1}}$, $\beta\geq 0$; (iv) substitute $\gamma$ inside the minimum operator. Moreover, (v) follows from the step size rule $\gamma=\frac{1}{8L_{1}(2\tau_{1}^{2}+1)(2\tau_{2}/(1-\tau_{1}))^{\beta}}\leq\frac{1}{4L_{1}(2\tau_{1}^{2}+1)(4\tau_{2}/(1-\tau_{1}))^{\beta}}$. Substituting equation 59 into equation 58 and summing over $t$ from 0 to T1T-1, we obtain that
@@
where (i) uses the concavity $\psi(\cdot)=y^{\frac{1}{2}}$ and $\psi(\cdot)=y^{\frac{1}{2-\beta}}$ and Jensen’s inequality $\psi(\frac{\sum_{i}x_{i}}{n})\geq\frac{\sum_{i}\psi(x_{i})}{n}$ applied on $\|\nabla F(w_{t})\|$; (ii) uses the fact $\mathbb{E}[\min\{A,B\}]\leq\mathbb{E}[A]\leq\max\{\mathbb{E}[A],\mathbb{E}[B]\}$; (iii) applies Jensen’s inequality over concave function $\Psi(\cdot)=y^{\frac{1}{2}}$ and $\Psi(\cdot)=y^{\frac{1}{2-\beta}}$; (iv) is due to re-organize equation 60, equation 61 and set K=TK={T}, $K^{c}=\frac{T}{2}$ (Since we require either $K$ or $K^{c}$ should be larger than $\frac{T}{2}$ to guarantee that $K+K_{c}=T$).
@@
$\overset{(i)}{=}$ $\gamma\min\Big\{1,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\gamma(\frac{1}{1-\tau_{1}}\big\|\nabla f_{\xi^{\prime}}(w_{t})\big\|+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\Big\}\big\|\nabla F(w_{t})\big\|^{2}$
@@
$\overset{(iv)}{=}$ $\min\Big\{\gamma,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\big(2C_{1}\big\|\nabla F(w_{t})\big\|\big)^{\beta}},\frac{1}{4L_{1}(2\tau_{1}^{2}+1)(2C_{2})^{\beta}}\Big\}\big\|\nabla F(w_{t})\big\|^{2}$
$\overset{(v)}{=}$ $\min\Big\{\gamma,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\big(2C_{1}\big\|\nabla F(w_{t})\big\|\big)^{\beta}}\Big\}\big\|\nabla F(w_{t})\big\|^{2}$